# The LHC has ruled out Supersymmetry – really?

### v1

@@ 1 Introduction
Supersymmetry (SUSY) is a theoretical framework that extends the Standard Model (SM) by a new symmetry between bosons and fermions. In this scenario, every known particle (more precisely, every SM degree of freedom) has a supersymmetric partner, differing in spin by half a unit. SUSY has been the prime paradigm for new physics beyond the SM for decades. For a detailed introduction, see e.g. Martin (1998).
One of the key motivations is the so-called hierarchy problem: within the SM, quantum corrections drive the Higgs boson mass to extremely high values, requiring a fine-tuning to maintain it at the electroweak scale. Weak-scale SUSY naturally solves this issue by ensuring that the divergent corrections from fermionic and bosonic partners cancel out, thereby stabilising the Higgs mass. Additionally, SUSY predicts the unification of gauge couplings at high energy scales. Another compelling aspect of supersymmetric models is the ability to provide viable dark matter candidates, particularly in models where R-parity is conserved: in this case, the lightest supersymmetric particle (LSP), often assumed to be a neutralino, remains stable and could account for the observed dark matter density in the universe.
@@
Early expectations for the discovery of SUSY at the LHC were high, particularly for strongly interacting superpartners. Theoretical arguments, but also global fits of constrained SUSY scenarios —such as the CMSSM— incorporating constraints from flavour physics, low-energy precision measurements, and dark matter relic density (see e.g. Allanach et al. (2007); Buchmueller et al. (2008); Trotta et al. (2008)), suggested that superparticle masses should be within reach of the LHC initial runs, typically below or around 1 TeV. However, searches by the ATLAS and CMS experiments (for a review, see e.g. Canepa (2019)) have ruled out substantial regions of this parameter space, challenging initial expectations. [^1] Comprehensive overviews of search results, including summary plots, are available at the ATLAS and CMS wiki pages https://twiki.cern.ch/twiki/bin/view/AtlasPublic/SearchesPublicResults and https://cms-results.web.cern.ch/cms-results/public-results/publications/SUS/index.html
Indeed, mass limits are being pushed higher and higher by the LHC experiments. The current bounds reported by ATLAS and CMS reach up to about $2.4$ TeV for gluinos and squarks of the first two generations, $1.3$ TeV for stops and sbottoms, and $1$ TeV for electroweak-inos [Kwon (2025)]. It has to be kept in mind, however, that these bounds are derived in the context of simplified models, assuming the presence of only few (typically 2–3) new particles and simple decay patterns (often a single decay channel). Realistic SUSY models can be much more complicated, with a large variety of relevant states and intricate mixing patterns affecting cross sections, decay branching ratios, and in specific cases even the lifetimes of the new states. Thus, even in the simplest realistic model, the so-called Mininal Supersymmetric Standard Model (MSSM), mass bounds can be weaker than those listed above.
@@ 2 Higgs mass, hierarchy and naturalness problems
where the cutoff $\Lambda$ represents the scale up to which the SM remains valid; this may be $M_{X}\geq 10^{15}$ GeV in a grand unified theory (GUT), or the Planck scale $M_{\rm P}\sim 10^{19}$ GeV, if no new physics sets in before. Moreover, in GUTs, the couplings of a light Higgs boson to the heavy Higgs fields present in the theory also yield $\delta m_{H}^{2}={\cal O}(M_{X}^{2})$. This is in principle not a problem in a renormalisable theory. However, it requires an extreme fine-tuning of counterterms order by order in pertubation theory, and this seems unnatural.
Figure 1: One-loop corrections to $m_{H}^{2}$, due to a Dirac fermion $f$ (left), and a scalar $S$ (right); from Martin (1998).
@@
$\delta m_{H}^{2}={\cal O}\left({\alpha\over 4\pi}\right)\;\left|m_{f}^{2}-m_{S}^{2}\,\right|\,,$ (2)
where $m_{f}$ and $m_{S}$ stand for the masses of the SM particles and their superpartners. This is no larger than the physical value, $\delta m_{H}^{2}\lesssim m_{H}^{2}$, provided
[…]
@@
Here $\mu$ is a SUSY-conserving Higgs mixing parameter from the superpotential (${\cal W}\in\mu H_{u}H_{d}$), while $m_{H_{u}}^{2}$, $m_{H_{d}}^{2}$ and $b$ are soft SUSY-breaking parameters; see Martin (1998) for details. The structure is that of a two-Higgs doublet model of type II and leads to five physical Higgs states, a neutral CP-odd, two neutral CP-even, and two charged Higgs bosons denoted by $A^{0}$, $h^{0}$, $H^{0}$, and $H^{\pm}$, respectively. An important point here is that the quartic term in eq. (4) is a function of the gauge couplings, in contrast to being a free parameter in the SM. As a consequence, after electroweak symmetry breaking, there is an upper bound on the mass of the lighter of the two CP-even Higgs bosons $h^{0}$:
@@
Figure 2: Dependence of the light MSSM Higgs boson mass, $M_{h}$, on the ratio $X_{t}/M_{S}$; from Arbey et al. (2012a).
@@
Here $M_{S}^{2}=m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}$ with $m_{\tilde{t}_{1,2}}$ the stop masses, and $X_{t}=A_{t}-\mu\cot\beta$ is the stop mixing parameter. To achieve the observed value of $m_{h}=125$ GeV [Aad et al. (2012); Chatrchyan et al. (2012)] either the average stop mass must be large, $M_{S}\equiv\sqrt{m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}}\gtrsim 3$ TeV, or the stop mixing parameter $|X_{t}|$ must be around twice $M_{S}$, as illustrated in Fig. 2 (taken from Arbey et al. (2012a), who, together with Brummer et al. (2012), also discuss the implications for GUT models). [^2] There is a vast literature on this subject, and on the topic of precision calculations of $m_{h}$. For reviews, we refer the reader to [Djouadi (2008); Draper and Rzehak (2016)]. This leads to a tension with the expectation of light stops (and not too heavy gluinos, due to their loop contributions to stop masses) motivated by the hierarchy problem, and with the wish for small radiative corrections to satisfy naturalness arguments.
@@
Models beyond the MSSM can alleviate the problem through extra tree-level contributions to $m_{h}^{2}$. This is the case, for instance, in models with extra singlet scalars, see e.g. [Ellwanger et al. (2010)], or when gauginos are Dirac instead of Majorana states [Benakli et al. (2013)]. Either way a tension remains and has led many people to doubt SUSY as a solution to the hierarchy and naturalness problems.
@@
One way out is to abandon the hierarchy problem altogether and use gauge-coupling unification and dark matter as the only guiding principles, as proposed in Arkani-Hamed and Dimopoulos (2005); Giudice and Romanino (2004); Arkani-Hamed et al. (2005). In this case, all the MSSM scalars become (ultra)heavy, except for a single finely tuned SM-like Higgs boson. Yet, the fermionic superpartners (gauginos and higgsinos) can remain light, protected by chiral symmetry, and account for the successful unification of gauge couplings. This is known as Split SUSY and has distinctive consequences for LHC signatures.
@@
An even more extreme approach is to consider that gauginos and higgsinos are also very heavy, with masses near the SUSY-breaking scale $M_{S}$. This defines the High-Scale SUSY framework Hall and Nomura (2010), where supersymmetry is effectively decoupled from the low-energy spectrum. In this scenario, one forgoes not only the naturalness argument but also the supersymmetric explanation for dark matter—since the LSP is no longer light—and the successful gauge coupling unification typically offered by SUSY. Nevertheless, supersymmetry leaves an imprint at low energies through the matching conditions between the high-scale SUSY theory and the effective Standard Model. In particular, the Higgs quartic coupling $\lambda$ retains information about the underlying SUSY structure. As a result, even when SUSY is broken at very high energies, it still predicts a relatively light Higgs boson, with its mass carrying indirect sensitivity to parameters like $M_{S}$ and $\tan\beta$.
@@
The relation between the observed Higgs mass at 125 GeV and the SUSY mass scale in the Split SUSY and High-Scale SUSY scenarios was investigated by Giudice and Strumia (2012); Arbey et al. (2012a); Bagnaschi et al. (2014) and is illustrated in Fig. 3.333This may be strongly influenced by tree-level contributions to $\lambda$ from a non-minimal field content.
@@
Figure 3: Predictions for the Higgs mass $m_{h}$ in Split SUSY (red) and in High-Scale SUSY (blue) for $\tan\beta=\{1,2,4,50\}$; from Giudice and Strumia (2012).
@@ 3 Little Hierarchy problem and electroweak naturalness
where the $\Sigma_{u,d}^{u,d}$ contain a large assortment of SUSY loop corrections. Barbieri and Giudice (1988) introduced the naturalness measure
@@
where the $p_{i}$ are fundamental parameters of the theory, and used this to set bounds on SUSY masses by imposing a maximally allowed value of $\Delta_{\rm BG}^{\rm max}=10$. Following this idea, Baer et al. (2016) formulated a naturalness principle stating that “An observable ${\cal O}$ is natural if all independent contributions to ${\cal O}$ are less than or of order ${\cal O}$”. This is dubbed “practical naturalness”. Applied to eq. (7), practical naturalness requires each term on the right-hand-side (RHS) of the equation to be $\lesssim m_{Z}^{2}/2$. This is used to define an electroweak fine-tuning measure $\Delta_{\rm EW}$ as
@@
Baer et al. (2016) conclude that gluinos, top squarks and other SUSY particles can live at the multi-TeV scale with little cost to naturalness (because they have loop-suppressed contributions to the weak scale), but higgsinos, whose masses are essentially given by $\mu$, must lie close to the weak scale, with $\mu\lesssim 200-400$ GeV. As we will see below, such light higgsinos have a tendency to escape detection at the LHC. Even at the High-Luminosity LHC, much but not all of the $\mu\lesssim 200-400$ GeV parameter space can be covered, see Baer et al. (2020).
@@
Figure 4: Range of sparticle and heavy Higgs masses allowed by naturalness within the NUHM2 model with $\Delta_{\rm EW}<30$. Also imposed are the LEP2 limit on charginos, LHC Run 1 bounds on gluino and squark masses, and $m_{h}=125\pm 2$ GeV. The black bars show upper bounds from the pMSSM model with 19 weak scale parameters; from Baer et al. (2016).
@@ 4 Phenomenological MSSM
The pMSSM is a 19-dimensional realisation of the R-parity conserving MSSM with parameters defined at the SUSY scale $M_{S}=\sqrt{m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}}$. It was introduced in Djouadi et al. (1998) to provide a more general framework for studying MSSM phenomenology than the predominant minimal or constrained SUSY-breaking models, without having to deal with the $>100$ free parameters of the general MSSM, which would be unwieldy for most phenomenological purposes.
Figure 5: Fraction of pMSSM points that are not excluded at 95% confidence level by ATLAS SUSY searches from Run 2 (given in Table 4 of Arbey et al. (2022)), and that remain beyond the projected sensitivity of Run 3, shown as a function of the gluino mass (upper panel) and the lighter stop mass (lower panel); from Arbey et al. (2022).
@@
This setup allows for generic mass and mixing patterns, leading to a much larger variety of cross sections, decay channels and branching ratios, and thus to a much broader set of predictions for experimental signatures than those found in any of the constrained scenarios, such as the CMSSM, mSUGRA, NUHM models, etc., which impose quite restrictive relations between the soft SUSY-breaking parameters (see AbdusSalam et al. (2010); Berger et al. (2009); Conley et al. (2011); Sekmen et al. (2012); Arbey et al. (2012b) for pioneering studies). Likewise, it also leads to a much broader set of predictions than realised in simplified models.
@@
Consequently, a significant portion of the pMSSM parameter space escapes the bounds from ATLAS and CMS searches, which were set within constrained or simplified models. This was confirmed by the ATLAS and CMS reinterpretations of their Run 1 searches, see Aad et al. (2015); Khachatryan et al. (2016). Indeed, mass limits derived within simplified models are often overly optimistic and do not hold in the more general case of the pMSSM (or other realistic SUSY models).
@@
As an example, Fig. 5 shows the fraction of pMSSM points (after flavour, dark matter, and low-energy constraints obtained using the SuperIso program Mahmoudi (2008, 2009)) not excluded by ATLAS Run 2 searches as a function of the gluino and stop masses. In the context of simplified models, the considered ATLAS analyses (given in Table 4 of Arbey et al. (2022)) exclude gluino masses up to $2.2-2.35$ TeV depending on the LSP mass, and stop masses in the range $450-950$ GeV for LSP masses below 160 GeV (in the case where $m_{\tilde{t}_{1}}\sim m_{t}+m_{\tilde{\chi}^{0}_{1}}$, stops in the range of $235-590$ GeV are excluded). This is to be compared with the full lines in Fig. 5, which indicate that gluinos with masses as low as 1 TeV and stops as light as 400 GeV remain viable within the pMSSM framework. [^4] It should be noted that the figure is not intended to provide a statistical interpretation; rather, it serves to illustrate that viable spectra with relatively light superpartner masses are indeed allowed within the pMSSM framework. Also shown in Fig. 5 are the projected sensitivities of Run 3 with 400/fb (dashed lines). As one can see, the fraction of surviving points is reduced, but the overall conclusion remains unchanged.
@@
Turning to electroweak-inos (i.e. charginos and neutralinos, mixtures of electroweak gauginos and higgsinos), in Aad et al. (2024) the ATLAS collaboration provided an interpretation of their Run 2 searches for electroweak production of supersymmetric particles within the pMSSM. To this end, they scanned bino, wino and higgsino mass parameters $M_{1}$, $M_{2}$ and $\mu$ from $-2$ TeV to $2$ TeV; the gluino mass parameter $M_{3}=1-5$ TeV; $m_{A}=0-5$ TeV and $\tan\beta=1-60$. The masses of 1st/2nd generation sfermions were fixed to 10 TeV and the 3rd generation parameters scanned over to obtain scenarios with a SM-like Higgs in the right mass range. Considering the results from eight separate searches for electroweak-inos (see Table 5 in Aad et al. (2024)), each using 140/fb of proton-proton data from Run 2, it was found that many pMSSM scenarios with chargino and neutralino masses well below the simplified-model limits escape exclusion, as examplified in Fig. 6.
@@
Figure 6: Fraction of electroweak-ino scenarios excluded by ATLAS Run 2 results. Considered are 7956 pMSSM points that have a neutralino LSP, satisfy chargino mass constraint from LEP and have a light Higgs mass in the range $m_{h}=120-130$ GeV. The white line shows the limit in the simplified wino/bino model; from Aad et al. (2024).
@@
Apparently, in the pMSSM, the total SUSY signal tends to spread out over several experimental channels. It is therefore interesting to ask how the sensitivity improves when statistically combining the likelihoods from different, independent analyses. Likewise, it is interesting to combine ATLAS and CMS results, to increase statistics and mitigate fluctuations in the data. This is possible using the SModelS framework as demonstrated recently in Altakach et al. (2023, 2024b). Since Aad et al. (2024) made all their scan points publicly available, we can use SModelS to reevaluate the constraints in a global likelihood analysis. Concretely, we are combining up to 33 analyses [^5] The best (i.e. most sensitive) combination is determined for each parameter point separately. Only analyses that can to good approximation be considered as uncorrelated enter the combination. from both Run 1 and Run 2 (21 from ATLAS and 12 from CMS) implemented in the SModelS v3.0.0 database [Altakach et al. (2024a)] that are sensitive to the electroweak-ino pMSSM scan points.
@@
The result is shown in Fig. 7. The plot shows (in colour) the allowed points, i.e. the points not excluded at 95% confidence level by the global likelihood, in the plane of lightest neutralino ($\tilde{\chi}^{0}_{1}$) versus lightest chargino ($\tilde{\chi}^{\pm}_{1}$) mass. The colour code indicates the combined observed $r$-value, $r^{\rm comb}_{\rm obs}$, with $r$ defined as the ratio of the predicted fiducial cross section of the signal over the corresponding upper limit. [^6] Points allowed by the combination but excluded in Aad et al. (2024) by the disappearing tracks analysis with 140/fb, for which only the 36/fb results are available in SModelS, have been removed. The grey points have $r^{\rm comb}_{\rm obs}\geq 1$ and are thus excluded by the combination.
@@
As one can see, only a small region of the pMSSM parameter space is definitely excluded, while many scenarios with light charginos and neutralinos remain valid. The colour code in Fig. 7 gives an idea about future prospects: very roughly, the red/orange/yellow-ish points should be within reach with more data at Run 3, while the blue(-ish) points will be hard to test (pending the excesses discussed in the next section).
@@
Figure 7: Scatter plot of pMSSM scan points from Aad et al. (2024) that escape exclusion in a global likelihood analysis with SModelS (in colour). Grey points are excluded by the combination of analyses. See text for details.

### v2

@@ 1 Introduction
Supersymmetry (SUSY) is a theoretical framework that extends the Standard Model (SM) by a new symmetry between bosons and fermions. In this scenario, every known particle (more precisely, every SM degree of freedom) has a supersymmetric partner, [^1] Here, we consider SUSY with $\mathcal{N}=1$ generator. For $\mathcal{N}>1$, supermultiplets grow larger. differing in spin by half a unit. SUSY has been the prime paradigm for new physics beyond the SM for decades. For a detailed introduction, see e.g. [1].
One of the key motivations is the so-called hierarchy problem: within the SM, quantum corrections drive the Higgs boson mass to extremely high values, requiring a fine-tuning to maintain it at the electroweak scale. Weak-scale SUSY naturally solves this issue by ensuring that the quadratically divergent corrections from fermionic and bosonic partners cancel out, thereby stabilising the Higgs mass. Moreover, through renormalization group running, the large top Yukawa coupling can drive the mass-squared term of (one of) the Higgs doublet scalars negative near the electroweak scale, thus naturally triggering electroweak symmetry breaking. Additionally, SUSY predicts the unification of gauge couplings at high energy scales. Another compelling aspect of supersymmetric models is the ability to provide viable dark matter candidates, particularly in models where R-parity is conserved: in this case, the lightest supersymmetric particle (LSP), often assumed to be a neutralino, remains stable and could account for the observed dark matter density in the universe.
@@
Early expectations for the discovery of SUSY at the LHC were high, particularly for strongly interacting superpartners. Theoretical arguments, but also global fits of constrained SUSY scenarios —such as the Constrained Mininal Supersymmetric Standard Model (CMSSM)— incorporating constraints from flavour physics, low-energy precision measurements, and dark matter relic density (see e.g. [2, 3, 4]), suggested that superparticle masses should be within reach of the LHC initial runs, typically below or around 1 TeV. However, searches by the ATLAS and CMS experiments (for a review, see e.g. [5]) have ruled out substantial regions of this parameter space, challenging initial expectations. [^2] Comprehensive overviews of search results, including summary plots, are available at the ATLAS and CMS wiki pages https://twiki.cern.ch/twiki/bin/view/AtlasPublic/SearchesPublicResults and https://cms-results.web.cern.ch/cms-results/public-results/publications/SUS/index.html
Indeed, mass limits are being pushed higher and higher by the LHC experiments. The current bounds reported by ATLAS and CMS reach up to about $2.4$ TeV for gluinos and squarks of the first two generations, $1.3$ TeV for stops and sbottoms, and $1$ TeV for electroweak-inos [6]. It has to be kept in mind, however, that these bounds are derived in the context of simplified models, assuming the presence of only few (typically 2–3) new particles and simple decay patterns (often a single decay channel). Realistic SUSY models can be much more complicated, with a large variety of relevant states and intricate mixing patterns affecting cross sections, decay branching ratios, and in specific cases even the lifetimes of the new states. Thus, even in the simplest realistic model, the so-called Mininal Supersymmetric Standard Model (MSSM), mass bounds can be weaker than those listed above.
@@ 2 Higgs mass, hierarchy and naturalness problems
where the cutoff $\Lambda$ represents the scale up to which the SM remains valid; this may be $M_{X}\geq 10^{15}$ GeV in a grand unified theory (GUT), or the Planck scale $M_{\rm P}\sim 10^{19}$ GeV, if no new physics sets in before. Moreover, in GUTs, the couplings of a light Higgs boson to the heavy Higgs fields present in the theory also yield $\delta m_{H}^{2}={\cal O}(M_{X}^{2})$. This is in principle not a problem in a renormalisable theory. However, it requires an extreme fine-tuning of counterterms order by order in pertubation theory, and this seems unnatural. [^3] For an in-depth discussion of the hierarchy problem, we refer to the article by M.E. Peskin [7] in this volume.
Figure 1: One-loop corrections to $m_{H}^{2}$, due to a Dirac fermion $f$ (left), and a scalar $S$ (right); from [1].
@@
$\delta m_{H}^{2}={\cal O}\left(\frac{\alpha}{4\pi}\right)\left(m_{S}^{2}-m_{f}^{2}\right)\log\left(\frac{\Lambda^{2}}{Q^{2}}\right)\,,$ (2)
where $m_{f}$ and $m_{S}$ stand for the masses of the SM particles and their superpartners, and $Q$ is typically of the order of SUSY-breaking scale. Thus, a residual logarithmic sensitivity to the ultraviolet cutoff remains due to soft SUSY breaking. This is no larger than the physical value, $\delta m_{H}^{2}\lesssim m_{H}^{2}$, provided
$|m_{S}^{2}-m_{f}^{2}|\lesssim\mathcal{O}(1~{}\text{TeV})^{2}\cdot\left[\log\left(\frac{\Lambda^{2}}{Q^{2}}\right)\right]^{-1}\,.$ (3)
@@
Here $\mu$ is a SUSY-conserving Higgs mixing parameter from the superpotential (${\cal W}\in\mu H_{u}H_{d}$), while $m_{H_{u}}^{2}$, $m_{H_{d}}^{2}$ and $b$ are soft SUSY-breaking parameters; see [1] for details. The structure is that of a two-Higgs doublet model of type II and leads to five physical Higgs states, a neutral CP-odd, two neutral CP-even, and two charged Higgs bosons denoted by $A^{0}$, $h^{0}$, $H^{0}$, and $H^{\pm}$, respectively. An important point here is that the quartic term in eq. (4) is a function of the gauge couplings, in contrast to being a free parameter in the SM. As a consequence, after electroweak symmetry breaking, there is an upper bound on the mass of the lighter of the two CP-even Higgs bosons $h^{0}$:
@@
Figure 2: Dependence of the light MSSM Higgs boson mass, $M_{h}$, on the ratio $X_{t}/M_{S}$; from [8].
@@
Here $M_{S}^{2}=m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}$ with $m_{\tilde{t}_{1,2}}$ the stop masses, and $X_{t}=A_{t}-\mu\cot\beta$ is the stop mixing parameter. To achieve the observed value of $m_{h}=125$ GeV [9, 10] either the average stop mass must be large, $M_{S}\equiv\sqrt{m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}}\gtrsim 3$ TeV, or the stop mixing parameter $|X_{t}|$ must be around twice $M_{S}$, as illustrated in Fig. 2 (taken from [8], who, together with [11], also discuss the implications for GUT models). [^4] There is a vast literature on this subject, and on the topic of precision calculations of $m_{h}$. For reviews, we refer the reader to [12, 13]. This leads to a tension with the expectation of light stops (and not too heavy gluinos, due to their loop contributions to stop masses) motivated by the hierarchy problem, and with the wish for small radiative corrections to satisfy naturalness arguments.
@@
Models beyond the MSSM can alleviate the problem through extra tree-level contributions to $m_{h}^{2}$. This is the case, for instance, in models with extra singlet scalars [14], or when gauginos are Dirac instead of Majorana states [15]. Either way a tension remains and has led many people to doubt SUSY as a solution to the hierarchy and naturalness problems.
@@
One way out is to abandon the hierarchy problem altogether and use gauge-coupling unification and dark matter as the only guiding principles, as proposed in [16, 17, 18]. In this case, all the MSSM scalars become (ultra)heavy, except for a single finely tuned SM-like Higgs boson. Yet, the fermionic superpartners (gauginos and higgsinos) can remain light, protected by chiral symmetry, and account for the successful unification of gauge couplings. This is known as Split SUSY and has distinctive consequences for LHC signatures.
@@
An even more extreme approach is to consider that gauginos and higgsinos are also very heavy, with masses near the SUSY-breaking scale $M_{S}$. This defines the High-Scale SUSY framework [19], where supersymmetry is effectively decoupled from the low-energy spectrum. In this scenario, one forgoes not only the naturalness argument but also the supersymmetric explanation for dark matter—since the LSP is no longer light—and the successful gauge coupling unification typically offered by SUSY. Nevertheless, supersymmetry leaves an imprint at low energies through the matching conditions between the high-scale SUSY theory and the effective Standard Model. In particular, the Higgs quartic coupling $\lambda$ retains information about the underlying SUSY structure. As a result, even when SUSY is broken at very high energies, it still predicts a relatively light Higgs boson, with its mass carrying indirect sensitivity to parameters like $M_{S}$ and $\tan\beta$.
@@
The relation between the observed Higgs mass at 125 GeV and the SUSY mass scale in the Split SUSY and High-Scale SUSY scenarios was investigated in [20, 8, 21] and is illustrated in Fig. 3.555This may be strongly influenced by tree-level contributions to $\lambda$ from a non-minimal field content.
@@
Figure 3: Predictions for the Higgs mass $m_{h}$ in Split SUSY (red) and in High-Scale SUSY (blue) for $\tan\beta=\{1,2,4,50\}$; from [20].
@@ 3 Little Hierarchy problem and electroweak naturalness
where the $\Sigma_{u,d}^{u,d}$ contain a large assortment of SUSY loop corrections. Barbieri and Giudice [22] introduced the naturalness measure
@@
where the $p_{i}$ are fundamental parameters of the theory, and used this to set bounds on SUSY masses by imposing a maximally allowed value of $\Delta_{\rm BG}^{\rm max}=10$. Following this idea, Ref. [23] formulated a naturalness principle stating that “An observable ${\cal O}$ is natural if all independent contributions to ${\cal O}$ are less than or of order ${\cal O}$”. This is dubbed “practical naturalness”. Applied to eq. (7), practical naturalness requires each term on the right-hand-side (RHS) of the equation to be $\lesssim m_{Z}^{2}/2$. This is used to define an electroweak fine-tuning measure $\Delta_{\rm EW}$ as
@@
The authors of [23] conclude that gluinos, top squarks and other SUSY particles can live at the multi-TeV scale with little cost to naturalness (because they have loop-suppressed contributions to the weak scale), but higgsinos, whose masses are essentially given by $\mu$, must lie close to the weak scale, with $\mu\lesssim 200-400$ GeV. As we will see below, such light higgsinos have a tendency to escape detection at the LHC. Even at the High-Luminosity LHC, much but not all of the $\mu\lesssim 200-400$ GeV parameter space can be covered, see [24].
@@
Figure 4: Range of sparticle and heavy Higgs masses allowed by naturalness within the NUHM2 model with $\Delta_{\rm EW}<30$. Also imposed are the LEP2 limit on charginos, LHC Run 1 bounds on gluino and squark masses, and $m_{h}=125\pm 2$ GeV. The black bars show upper bounds from the pMSSM model with 19 weak scale parameters; from [23].
@@ 4 Phenomenological MSSM
The pMSSM is a 19-dimensional realisation of the R-parity conserving MSSM with parameters defined at the SUSY scale $M_{S}=\sqrt{m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}}$. It was introduced in [25] to provide a more general framework for studying MSSM phenomenology than the predominant minimal or constrained SUSY-breaking models, without having to deal with the $>100$ free parameters of the general MSSM, which would be unwieldy for most phenomenological purposes.
Figure 5: Fraction of pMSSM points that are not excluded at 95% confidence level by ATLAS SUSY searches from Run 2 (given in Table 4 of [26]), and that remain beyond the projected sensitivity of Run 3, shown as a function of the gluino mass (upper panel) and the lighter stop mass (lower panel); from [26].
@@
This setup allows for generic mass and mixing patterns, leading to a much larger variety of cross sections, decay channels and branching ratios, and thus to a much broader set of predictions for experimental signatures than those found in any of the constrained scenarios, such as the CMSSM, mSUGRA666minimal SUperGRAvity (mSUGRA) is a scenario in which SUSY breaking is mediated by gravity., NUHM models, etc., which impose quite restrictive relations between the soft SUSY-breaking parameters (see [27, 28, 29, 30, 31] for pioneering studies). Likewise, it also leads to a much broader set of predictions than realised in simplified models.
@@
Consequently, a significant portion of the pMSSM parameter space escapes the bounds from ATLAS and CMS searches, which were set within constrained or simplified models. This was confirmed by the ATLAS and CMS reinterpretations of their Run 1 searches [32, 33]. Indeed, mass limits derived within simplified models are often overly optimistic and do not hold in the more general case of the pMSSM (or other realistic SUSY models).
@@
As an example, Fig. 5 shows the fraction of pMSSM points (after flavour, dark matter, and low-energy constraints obtained using the SuperIso program [34, 35]) not excluded by ATLAS Run 2 searches as a function of the gluino and stop masses. In the context of simplified models, the considered ATLAS analyses (given in Table 4 of [26]) exclude gluino masses up to $2.2-2.35$ TeV depending on the LSP mass, and stop masses in the range $450-950$ GeV for LSP masses below 160 GeV (in the case where $m_{\tilde{t}_{1}}\sim m_{t}+m_{\tilde{\chi}^{0}_{1}}$, stops in the range of $235-590$ GeV are excluded). This is to be compared with the full lines in Fig. 5, which indicate that gluinos with masses as low as 1 TeV and stops as light as 400 GeV remain viable within the pMSSM framework. [^7] It should be noted that the figure is not intended to provide a statistical interpretation; rather, it serves to illustrate that viable spectra with relatively light superpartner masses are indeed allowed within the pMSSM framework. Also shown in Fig. 5 are the projected sensitivities of Run 3 with 400/fb (dashed lines). As one can see, the fraction of surviving points is reduced, but the overall conclusion remains unchanged.
@@
Turning to electroweak-inos (i.e. charginos and neutralinos, mixtures of electroweak gauginos and higgsinos), in [36] the ATLAS collaboration provided an interpretation of their Run 2 searches for electroweak production of supersymmetric particles within the pMSSM. To this end, they scanned bino, wino and higgsino mass parameters $M_{1}$, $M_{2}$ and $\mu$ from $-2$ TeV to $2$ TeV; the gluino mass parameter $M_{3}=1-5$ TeV; $m_{A}=0-5$ TeV and $\tan\beta=1-60$. The masses of 1st/2nd generation sfermions were fixed to 10 TeV and the 3rd generation parameters scanned over to obtain scenarios with a SM-like Higgs in the right mass range. Considering the results from eight separate searches for electroweak-inos (see Table 5 in [36]), each using 140/fb of proton-proton data from Run 2, it was found that many pMSSM scenarios with chargino and neutralino masses well below the simplified-model limits escape exclusion, as exemplified in Fig. 6.
@@
Figure 6: Fraction of electroweak-ino scenarios excluded by ATLAS Run 2 results. Considered are 7956 pMSSM points that have a neutralino LSP, satisfy chargino mass constraint from LEP and have a light Higgs mass in the range $m_{h}=120-130$ GeV. The white line shows the limit in the simplified wino/bino model; from [36].
@@
Apparently, in the pMSSM, the total SUSY signal tends to spread out over several experimental channels. It is therefore interesting to ask how the sensitivity improves when statistically combining the likelihoods from different, independent analyses. Likewise, it is interesting to combine ATLAS and CMS results, to increase statistics and mitigate fluctuations in the data. This is possible using the SModelS framework as demonstrated recently in [37, 38]. Since [36] made all their scan points publicly available, we can use SModelS to reevaluate the constraints in a global likelihood analysis. Concretely, we are combining up to 33 analyses from both Run 1 and Run 2 (21 from ATLAS and 12 from CMS) implemented in the SModelS v3.0.0 database [39] that are sensitive to the electroweak-ino pMSSM scan points. The best (i.e. most sensitive) combination is determined for each parameter point separately. Only analyses that can to good approximation be considered as uncorrelated enter the combination.
@@
The result is shown in Fig. 7. The plot shows (in colour) the allowed points, i.e. the points not excluded at 95% confidence level by the global likelihood, in the plane of lightest neutralino ($\tilde{\chi}^{0}_{1}$) versus lightest chargino ($\tilde{\chi}^{\pm}_{1}$) mass. The colour code indicates the combined observed $r$-value, $r^{\rm comb}_{\rm obs}$, with $r$ defined as the ratio of the predicted fiducial cross section of the signal over the corresponding upper limit. [^8] Points allowed by the combination but excluded in [36] by the disappearing tracks analysis with 140/fb, for which only the 36/fb results are available in SModelS, have been removed. The grey points have $r^{\rm comb}_{\rm obs}\geq 1$ and are thus excluded by the combination.
@@
As one can see, only a small region of the pMSSM parameter space is definitely excluded, while many scenarios with light charginos and neutralinos remain valid. [^9] A similar conclusion was reached in [40], studying the validity of light neutralino thermal dark matter. The colour code in Fig. 7 gives an idea about future prospects: very roughly, the red/orange/yellow-ish points should be within reach with the projected 500 fb-1 of data at Run 3, while the blue(-ish) points may be hard to test (pending the excesses discussed in the next section).
@@
Figure 7: Scatter plot of pMSSM scan points from [36] that escape exclusion in a global likelihood analysis with SModelS (in colour). Grey points are excluded by the combination of analyses. See text for details.